# LLM-Generated Black-box Explanations Can Be Adversarially Helpful

### v2

@@ Abstract
Large Language Models (LLMs) are becoming vital tools that help us solve and understand complex problems by acting as digital assistants. LLMs can generate convincing explanations, even when only given the inputs and outputs of these problems, i.e., in a “black-box” approach. However, our research uncovers a hidden risk tied to this approach, which we call adversarial helpfulness. This happens when an LLM’s explanations make a wrong answer look right, potentially leading people to trust incorrect solutions. In this paper, we show that this issue affects not just humans, but also LLM evaluators. Digging deeper, we identify and examine key persuasive strategies employed by LLMs. Our findings reveal that these models employ strategies such as reframing the questions, expressing an elevated level of confidence, and cherry-picking evidence to paint misleading answers in a credible light. To examine if LLMs are able to navigate complex-structured knowledge when generating adversarially helpful explanations, we create a special task based on navigating through graphs. Most LLMs are not able to find alternative paths along simple graphs, indicating that their misleading explanations aren’t produced by only logical deductions using complex knowledge. These findings shed light on the limitations of the black-box explanation setting and allow us to provide advice on the safe usage of LLMs.
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong explaining capabilities. They can generate texts containing logical steps towards solving complex problems (Trinh et al., 2024; Sprague et al., 2023), and incorporate users’ situations (Mondal et al., 2024; Zhu et al., 2023; Zhou et al., 2024). The literature in NLP has many works showing LLMs can generate explanations that are convincing (Wiegreffe et al., 2022), informative (Chen et al., 2023a) and faithful (Lyu et al., 2023) to levels that are comparable to human-written explanations. These abilities lead to a promising use case: LLMs as assistants that explain problems to us.
Sounds like we have a convenient use case: a user can pass a complex problem and its answer to an LLM, together with a suitably-formatted prompt. Hopefully, this “black-box” approach will let the LLM generate an explanation that is instructive and helpful for us to understand the problem. Hopefully, this use case would facilitate education, understanding, and decision-making in the ocean of complex data. However, as we will show in this paper, this approach may backfire. Specifically, the explanations generated by this “black-box” approach may encourage humans to believe in the incorrect explanandum. We refer to this problem adversarial helpfulness.
We first try to gauge the extent of this adversarial helpfulness problem. We let human annotators rate the extent they are convinced of an incorrect answer, before and after seeing an LLM-generated explanation. On a commonsense reasoning task, their ratings are significantly improved by explanations generated by three commercial-tier LLMs. We also use three top-ranking LLMs as proxy evaluators to repeat the survey, and observe similar effects.
@@
Then, we proceed with studying the strategies of the adversarially helpful explanations. We identify ten strategies that we consider relevant to the explanations being adversarially helpful, and examine the strategies used by LLMs in generating these explanations. We find alarmingly high frequencies. For example, over 90% of the explanations in inference problems involve reframing to varying extents, and over 60% of the explanations in the commonsense problems involve the selective presentation of either the facts or the evidence.
Then we analyze the explanations from a reason graph perspective. Are these LLMs able to generate these adversarially helpful explanations because they are able to navigate through complex knowledge, like finding an alternate path in a graph? We set up a symbolic reasoning task, where we ask the LLMs to figure out an alternate path that leads to a specified destination node. This is an abstraction for “explaining an incorrect answer”. We find that the weaker models are unable to complete this task, especially when the graph complexity increases. These indicate that the generation of adversarially helpfulness explanations may involve more than the abilities in deductive reasoning and logical inference, which matches the prior observation that additional strategies (e.g., reframing) are used.
The findings shed light on the limitations of the black-box explanation setting, and motivate us to provide recommendations for safer usage of LLMs as explaining assistants. The access to all analysis scripts and data is open at https://github.com/ZiningZhu/adversarial_helpfulness.
@@ 2 Related Works
Recently, a number of works have leveraged the reasoning abilities of LLMs: chain-of-thought (Nye et al., 2021; Wei et al., 2022), tree-of-thought (Yao et al., 2024), graph-of-thought (Besta et al., 2023) and everything-of-thought (Ding et al., 2023) are representative examples. We follow their approaches to elicit the reasoning abilities from LLMs.
@@
Model-generated explanations can have significant impacts on both human users, for example in mitigating misinformation (Hsu et al., 2023; Si et al., 2024), rescaling human judgments (Wadhwa et al., 2023) and understanding models’ behavior (Hase & Bansal, 2020; Chen et al., 2022). Researchers have tried to use explanations for model development, yet there are mixed results (Saha et al., 2024; Im et al., 2023).
@@
The explainability research of AI has identified pitfalls (those unintended effects) versus dark patterns (those resulting from intended misuse) (Ehsan & Riedl, 2021). The adversarial helpfulness of LLM-generated explanation is an example of explainability pitfalls. The explanations towards models have been known to lead to problems including over-trust (Jacovi et al., 2021), over-reliance (Chen et al., 2023b) and incorrect calibrations (Zhang et al., 2020). Specific to the explanations in natural language, these explanations are selective and can be subjective or misleading (Kunz & Kuhlmann, 2024; Xu et al., 2023b) or unreliable (Ye & Durrett, 2022). Adversarial helpfulness is a different problem, as we’ll elaborate in the discussion section. LLM-generated natural language explanations about commonsense problems are perceived to have comparable attributes (generality, factuality, grammatical correctness, informativeness, acceptability) regardless of the correctness of the problems (Wiegreffe et al., 2022); we focus on the cases where the incorrect problems are explained in a post-hoc black-box manner.
@@
LLMs with self-explanations can be biased towards incorrect answers by superficial patterns like the ordering of choices (Turpin et al., 2024). Our paper considers a different setting: instead of planting superficial patterns in the demonstrations, we instruct the LLMs to explain an incorrect answer in a zero-shot manner, resembling how a user would use the LLM as an explaining assistant. This paper is also relevant to the jailbreaking/red-teaming of LLMs (Zou et al., 2023; Zeng et al., 2024; Deng et al., 2023; Ganguli et al., 2022). Instead of developing the jailbreaking or defending algorithms, we focus on the problem itself, and study the strategies adopted by the LLMs when generating adversarially helpful explanations.
@@ 3 Experiment setup
Two datasets are used: ECQA (Aggarwal et al., 2021a) and SNLI (Bowman et al., 2015). For ECQA, we sampled 500 problems and hand-pick a “second-best answer” that we consider to be only slightly worse than the correct answer designated by the dataset. Anecdotally, we found that those more abstract problems lead to more significant “adversarial helpfulness” explanations, so we skipped some concrete problems. For SNLI, we sampled 300 problems from each of the datasets with label Entailment and Contradictory, respectively. We found it was almost impossible to write sufficiently logical arguements (with humans or LLMs) that an Entailment sentence pair should be Contradictory (or vice versa), so we only considered the cases that argue for a Neutral label.
@@
We consider the following four models as explainers: ChatGPT (gpt-3.5-turbo) (OpenAI, 2022), GPT-4 (gpt-4-0613) (Achiam et al., 2023), Claude (Amazon Bedrock version) (Anthropic, 2023), Cohere Command (Cohere, 2023). These are powerful models that provide API access for text generation. We include only the API access models here, since the API models have a potentially larger audience who may be vulnerable to adversarially helpful explanations.
@@
We consider the following models as evaluators: Vicuna-33B-v1.3 (Chiang et al., 2023), WizardLM-70B-V1.0 (Xu et al., 2023a), Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). These are the top-ranking [^1] As of February 2024. Note that Qwen1.5-72B-Chat (Bai et al., 2023) and Yi-34B-Chat (AI et al., 2024) are also ranked on the top places, but we only included one LLM per model size to reduce the computation cost. models on the lmsys Chatbot Arena leaderboard (Zheng et al., 2023) that provide weight access. The weight access is important, since it allows us to query the probabilities in setting up the automatic evaluation — we will describe the setup in Section 5.
@@
The “adversarially helpful” explanations are generated by prompting the LLM explainer to explain an incorrect answer to a question. Figure 2 lists the prompting templates we used. In our preliminary experiments, we found other prompting templates can lead to explanations demonstrating a similar “adversarially helpful” behavior, as long as we specifically instruct the model to convince the audience of the incorrect answer.
@@ 4 Human evaluation of adversarial helpfulness
We recruit Amazon Mturk annotators to evaluate the following four scores, on GPT4-generated explanations for the ECQA problems. Each explanation is presented to three randomly recruited annotators, who selects from 1, 3 and 5 for each of the following four scores.
@@
Appendix A includes the screenshots of the UI, including the marking criteria. This protocol is approved by the ethical review board at our university.
@@
The MTurk annotators rate the LLM-generated explanations to have high fluency and correctness ratings. As Table 1 shows, the convincingness ratings rise from $2.96(\textrm{sd}=0.99)$ to $3.53(\textrm{sd}=0.93)$, from $3.66(\textrm{sd}=0.88)$ to $3.72(\textrm{sd}=0.85)$ and from $3.74(\textrm{sd}=0.97)$ to $3.84(\textrm{sd}=0.94)$, for GPT4, Claude, and ChatGPT, respectively. Paired $t$-tests (dof=$499$, two-tailed for all three) find significant differences ($p<0.01$ for all three) between the pre-explanation and the post-explanation convincingness scores, showing that the humans consider the explanations beneficial for the convincingness of the answers, even when the answers are incorrect.
5 Automatic evaluation of adversarial helpfulness
Evaluator setup
To examine the effects of the generated explanations in a scalable manner, we use several evaluator language models as proxies for MTurk annotators. Following is the protocol we use for querying the response from a proxy.
$\hat{y}=argmax_{y\in[\textrm{``1'',``3'',``5''}]}P(y\,|\,\mathbf{x})$ (1)
Here, the input texts $\mathbf{x}$ for each question are identical to the texts presented to the human annotators in MTurk. We observe several interesting effects, and summarize them below.
@@ 5 Automatic evaluation of adversarial helpfulness
We observe relatively consistent trends on both ECQA and NLI datasets. On both datasets, the convincingness ratings for the incorrect answers by Mixtral-8x7B do not significantly differ. The other two models, Vicuna-33B and WizardLM-70B, compute increased probabilities which show statistical significance ($p<0.001$ on two-tailed $t$ tests, Bonferroni corrected, with $\textrm{dof}=499$ for ECQA and $\textrm{dof}=299$ for NLI).
@@
Note that the utility of LLM that evaluates the convincingness should be taken with caution. First, LLMs have demonstrated evidence to model humans’ thoughts (i.e., “theory-of-mind” modeling) (Kosinski, 2024), but the actual capability is being debated. Second, in several cases, the LLM-given ratings show “degeneration” trends. For example, in NLI $E\rightarrow N$, all C_after scores computed by Mixtral-8x7B are the same, resulting in an average score of 3.00. This indicates that the scores given by the proxy evaluators might be affected by the dataset artifacts, in addition to the contents.
Models and humans show varying numbers, item-wise
When getting down to the item-wise level, the evaluators show varying trends. First, humans correlate weakly with the proxy models. On the dataset where human annotator results are available, we compute the Pearson correlations between the averaged human results and the evaluator models. None of the correlations are significant, indicating that the evaluator models show very different fluency, correctness and convincingness ratings from the human annotators. Second, human results show poor inter-annotator agreements. This is because the MTurk platform distributes the annotation tasks to more than three annotators. Third, the proxy models assign different scores. We compute the Cohen’s Kappa between each pair of the proxy models. None of them have a value larger than 1 for any score, indicating poor agreement between the evaluators. This is because the proxy models have different “baseline marking guidelines”, as is illustrated by the drastically different mean scores in Table 1.
6 Strategies toward adversarial helpfulness
The recent literature involves many taxonomies of persuasion strategies. For example, Dimitrov et al. (2021) identified strategies in social media texts. Piskorski et al. (2023) considered news. Zeng et al. (2024) considered 40 techniques for jailbreaking LLMs as well as other NLP tasks. Inspired by these works, we identify ten persuasion strategies that we consider particularly relevant to LLM-generated explanations. A brief summarization of each strategy is included in Appendix B.
We use GPT-4 Turbo (gpt-4-0125-preview) to detect the persuasion strategies used in the explanations. Appendix B lists the prompt. If no strategy is detected, the json object parsed from the LLM’s response would contain zero in all ten entries. Figure 1 lists two examples of the identified strategies.
Frequencies of the strategies
First, the sheer numbers of frequency are alarming. For the commonsense questions, more than $70\%$ of the explanations highlight the comparative advantage towards the incorrect answers. For inference tasks, more than $90\%$ of the explanations involve reframing of the questions. While these results are only on the commonsense and inference datasets, we expect that the real-world applications of natural language explanations frequently include these persuasion strategies. Do LLM explanations exhibit these strategies for correct problems? We hypothesize an affirmative answer, since LLMs have these persuasion capabilities. However, we argue that a safe explainer should utilize few or none of these persuasion strategies in explanation, especially when the explanandum involves an incorrect problem.
Second, the three explainer models have common trends in applying persuasion strategies. The LLM-generated explanations appear overly confident. [^2] This may related to the overly-confident tone in our prompts, as listed in Figure 2. Strategies like selective evidence, and selective fact presentation are frequently used. [^3] Relatedly, selectivity is a desirable feature in human explanations (Lombrozo & Liquin, 2023). The strategies like appeal to authority and analogical evidence are very infrequently used in any of the models. These indicate that safeguarding against adversarially helpful explanations might benefit from the identification and prevention of a finite set of persuasion strategies.
For completeness, we repeat the automatic detection experiments using the taxonomy of Zeng et al. (2024) — the experiment results are included in Table 3 in Appendix, showing similar observations: Only a few of the persuasion strategies/techniques are applied (e.g., logical appeal, encouragement and framing), but for those applied techniques, the frequencies are alarmingly high.
Future Works
This section presents an exploratory analysis of the explanation strategies, opening up future research directions. First, the cause of each strategy can be analyzed, by for example correlating each persuasion strategy and linguistic markers like syntactic complexity. Second, how each of the strategies affects the adversarial helpfulness can be studied in future works. Third, how each of the persuasion strategies can be mitigated, either by prompting techniques or by model alignment, would be promising future works.

### v3

@@ Abstract
Large language models (LLMs) are becoming vital tools that help us solve and understand complex problems. LLMs can generate convincing explanations, even when given only the inputs and outputs of these problems, i.e., in a “black-box” approach. However, our research uncovers a hidden risk tied to this approach, which we call adversarial helpfulness. This happens when an LLM’s explanations make a wrong answer look correct, potentially leading people to trust faulty solutions. In this paper, we show that this issue affects not just humans, but also LLM evaluators. Digging deeper, we identify and examine key persuasive strategies employed by LLMs. Our findings reveal that these models employ strategies such as reframing questions, expressing an elevated level of confidence, and ‘cherry-picking’ evidence that supports incorrect answers. We further create a symbolic graph reasoning task to analyze the mechanisms of LLMs generating adversarial helpfulness explanations. Most LLMs are not able to find alternative paths along simple graphs, indicating that other mechanisms, rather than logical deductions, might facilitate adversarial helpfulness. These findings shed light on the limitations of black-box explanations and lead to recommendations for the safer use of LLMs.
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong capabilities in explanation, including providing logical steps towards solving complex problems (Trinh et al., 2024; Sprague et al., 2023), incorporating user contexts (Mondal et al., 2024; Zhu et al., 2023; Zhou et al., 2024), and generating explanations that are convincing (Wiegreffe et al., 2022), informative (Chen et al., 2023a), and faithful (Lyu et al., 2023) to levels that are comparable to humans. These abilities lead to an apparently promising use case: LLMs as explainer assistants.
Under this use case, a user can pass a complex problem and its answer to an LLM, together with a suitably-formatted prompt. Hypothetically, this “black-box” approach lets the LLM generate an explanation that is instructive and helpful for us to understand the problem. Also hypothetically, this use case would facilitate education, understanding, and decision-making in an ocean of complex data. However, this approach may backfire. Specifically, the explanations generated by this “black-box” approach may encourage humans to believe in incorrect phenomena. We refer to this problem as adversarial helpfulness.
We first try to gauge the extent of this adversarial helpfulness problem. We let human annotators rate the extent they are convinced of an incorrect answer, before and after seeing an LLM-generated explanation. On a commonsense reasoning task, their ratings are significantly improved by explanations generated by three commercial-tier LLMs. We also use three top-ranking LLMs as proxy evaluators to repeat the survey and observe similar effects.
@@
We then study the strategies used in the adversarially helpful explanations. We identify ten strategies that we consider relevant to the explanations being adversarially helpful and examine the strategies used by LLMs in generating these explanations. We detect these strategies at alarmingly high frequencies. For example, over 90% of the explanations in inference problems involve reframing to varying extents, and over 60% of the explanations in the commonsense problems involve the selective presentation of either the facts or the evidence.
We analyze the explanations from a ‘reason graph’ perspective. Are these LLMs able to generate these adversarially helpful explanations because they can navigate through complex knowledge, like finding an alternate path in a graph? We set up a symbolic reasoning task, where we ask the LLMs to find an alternate path that leads to a specified destination “reasoning node”. We consider this to be an abstraction for “explaining an incorrect answer”. We find that the weaker models are unable to complete this task, especially when the graph complexity increases. These findings indicate that the generation of adversarially helpful explanations may involve more than the abilities in deductive reasoning and logical inference, which matches the prior observation that additional strategies (e.g., reframing) are used.
We shed light on the limitations of black-box explanation settings and provide recommendations for the safer use of LLMs as explaining assistants, towards ensuring the rights to explanations. Access to all analysis code and data is open at GitHub.
@@ 2 Related Works
Recent work has leveraged the reasoning abilities of LLMs, e.g., chain-of-thought (Nye et al., 2021; Wei et al., 2022), tree-of-thought (Yao et al., 2024), graph-of-thought (Besta et al., 2023) and everything-of-thought (Ding et al., 2023) are representative. We follow these approaches and leverage the reasoning abilities of LLMs.
@@
Model-generated explanations can have significant impacts on both human users, e.g., in answering the question (Joshi et al., 2023), mitigating misinformation (Hsu et al., 2023; Si et al., 2024), rescaling human judgments (Wadhwa et al., 2023), and understanding model behavior (Hase and Bansal, 2020; Chen et al., 2022). Researchers have tried to use explanations for model development, to mixed results (Saha et al., 2024; Im et al., 2023). Non-LLM explanation methods like feature contribution, gradient attribution, and input highlighting have also produced mixed results (Buçinca et al., 2020; Bansal et al., 2021; Wang and Yin, 2021; Kim et al., 2022).
@@
The explainability research in AI differentiates between ‘pitfalls’ (unintended effects) and ‘dark patterns’ (intended misuse) (Ehsan and Riedl, 2021). The adversarial helpfulness of LLM-generated explanations is a pitfall of explainability. Other pitfalls of explanations include over-trust (Jacovi et al., 2021), over-reliance (Chen et al., 2023b), and incorrect calibrations (Zhang et al., 2020). Researchers have also doubted the explanations of other prediction models in a post-hoc manner (Kroeger et al., 2023). Specific to the explanations in natural language, these explanations are selective and can be subjective, misleading (Kunz and Kuhlmann, 2024; Xu et al., 2023a), or unreliable (Ye and Durrett, 2022). Adversarial helpfulness is a different problem, as we’ll elaborate in Section 8.
LLM-generated natural language explanations about commonsense questions and answers are perceived to have comparable attributes (generality, factuality, grammatical correctness, informativeness, acceptability) to human-written explanations, regardless of the correctness of the answers (Wiegreffe et al., 2022); we focus on the cases where the incorrect problems are explained in a post-hoc black-box manner, and analyze the mechanisms of these explanations.
@@
LLMs with self-explanations can be biased towards incorrect answers by superficial patterns like the ordering of choices (Turpin et al., 2024). We consider a different setting: instead of planting superficial patterns in the demonstrations, we instruct the LLMs to explain an incorrect answer in a zero-shot manner, resembling how a user would use the LLM as an explaining assistant. This paper is relevant to the jailbreaking and red-teaming of LLMs (Zou et al., 2023; Zeng et al., 2024; Deng et al., 2023; Ganguli et al., 2022). Instead of developing jailbreaking or defense algorithms, we focus on the problem itself. We study the strategies adopted by the LLMs when generating adversarially helpful explanations and recommend targeted mitigation guidelines.
@@ 3 Experiment setup
Two datasets are used: ECQA (Aggarwal et al., 2021a) and SNLI (Bowman et al., 2015). The ECQA (Explanation-Centered Question Answering) dataset is designed to evaluate the quality of explanations provided by models, focusing on the clarity, relevance, and coherence of the generated explanations. For ECQA, we sample 500 problems and select a "second-best answer" that we consider to be only slightly worse than the correct answer designated by the dataset. Anecdotally, those more abstract problems lead to more significant "adversarial helpfulness" explanations. We are also not interested in the types of problems that require direct contradictions to the given facts, so we skip some concrete problems.
The SNLI (Stanford Natural Language Inference) dataset, on the other hand, is a large-scale collection of sentence pairs with labels indicating entailment, contradiction, or neutrality. It is widely used for training and evaluating models on the task of natural language inference. For SNLI, we sample 300 problems from each of the datasets with the Entailment label and the Contradictory label, respectively. We find that it is almost impossible to write sufficiently logical arguments (for humans or LLMs) to explain an Entailment sentence pair into a Contradictory pair (or vice versa), so we only consider the cases that explain for a Neutral label.
@@
We consider the following four models as explainers: Chat-3.5-Turbo (OpenAI, 2022), GPT-4 (gpt-4-0613) (Achiam et al., 2023), Claude (Amazon Bedrock version) (Anthropic, 2023), and Cohere Command (Cohere, 2023). These are powerful models that provide API access for text generation. We include only the API access models here because, compared to the weight-access models, the API-access models have a larger audience who may be vulnerable to adversarially helpful explanations.
@@
We consider the following models as evaluators: Vicuna-33B-v1.3 (Chiang et al., 2023), WizardLM-70B-V1.0 (Xu et al., 2023b), and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). These are the top-ranking [^1] As of February 2024. Note that Qwen1.5-72B-Chat (Bai et al., 2023) and Yi-34B-Chat (AI et al., 2024) are also top-ranked, but we only include one LLM per model size for feasibility. models on the lmsys Chatbot Arena leaderboard (Zheng et al., 2023) that provide weight access. The weight access is important since it allows us to query the probabilities in setting up the automatic evaluation — we describe the setup in Section 5.
@@
The “adversarially helpful” explanations are generated by prompting the LLM explainer to explain an incorrect answer to a question. Figure 2 lists the prompting templates we use. In our preliminary experiments, we found other prompting templates can lead to explanations demonstrating a similar “adversarially helpful” behavior, as long as we specifically instruct the model to convince the audience of the incorrect answer.
@@
Hypothesis: {{hypothesis}}
Answer choices: Entailment, Neutral, Contradiction
Convince the audience why {{incorrect-answer}} is definitely correct and why the other options are wrong. Write 3-4 lines for your explanation.
@@ 4 Human evaluation of adversarial helpfulness
We recruit Amazon Mturk annotators (from the US region) to evaluate the following four scores, on GPT 4-generated explanations for the ECQA problems. Each explanation is presented to three randomly recruited annotators, who select from 1, 3, and 5 for each of the following four scores. This configuration resembles a 3-point Likert scale.
@@
Appendix B includes screenshots of the UI, including the marking criteria. This protocol is approved by the ethical review board at our university.
@@
The MTurk annotators rate the LLM-generated explanations to have high fluency and correctness ratings. As Table 1 shows, the convincingness ratings rise from $2.96(\textrm{sd}=0.99)$ to $3.53(\textrm{sd}=0.93)$, from $3.66(\textrm{sd}=0.88)$ to $3.72(\textrm{sd}=0.85)$ and from $3.74(\textrm{sd}=0.97)$ to $3.84(\textrm{sd}=0.94)$, for GPT4, Claude, and GPT-3.5-Turbo, respectively. Paired $t$-tests (dof=$499$, two-tailed for all three) find significant differences ($p<0.01$ for all three) between the pre-explanation and the post-explanation convincingness scores, showing that the humans consider the explanations beneficial for the convincingness of the answers, even when the answers are incorrect.
@@
5 Automatic evaluation of adversarial helpfulness
Evaluator setup
To examine the effects of the generated explanations in a scalable manner, we use several evaluator language models as proxies for MTurk annotators. The following is the protocol we use for querying the response from a proxy.
$\hat{y}=argmax_{y\in[\textrm{``1'',``3'',``5''}]}P(y\,|\,\mathbf{x})$ (1)
Here, the input texts $\mathbf{x}$ for each question are identical to the texts presented to the human annotators in MTurk. We observe several interesting effects and summarize them below.
@@ 5 Automatic evaluation of adversarial helpfulness
We observe relatively consistent trends on both ECQA and NLI datasets. On both datasets, the convincingness ratings for the incorrect answers by Mixtral-8x7B do not significantly differ. The other two models, Vicuna-33B and WizardLM-70B, compute increased probabilities which show statistical significance ($p<0.001$ on two-tailed $t$-tests, Bonferroni corrected, with $\textrm{dof}=499$ for ECQA and $\textrm{dof}=299$ for NLI).
@@
Note that the utility of an LLM that evaluates the convincingness should be treated with caution. First, LLMs have demonstrated evidence of modeling human thoughts (i.e., “theory-of-mind” modeling) (Kosinski, 2024), but the actual capability is being debated. Second, in several cases, the ratings provided by LLMs show “degeneration” trends. For example, in NLI $E\rightarrow N$, all C_after scores computed by Mixtral-8x7B are identical (averaging 3.00). This indicates that the scores given by the proxy evaluators might be affected by the dataset artifacts, in addition to the contents.